# Emojis Decoded: Leveraging ChatGPT for Enhanced Understanding in Social Media Communications

### 22 Jan 2024 v1

@@ 1 Introduction
Large language models (LLMs), pre-training on large corpuses, have demonstrated the remarkable achievements in multiple natural language processing tasks. Among various LLMs, ChatGPT (GPT3.5, GPT4, and GPT4V) (Ouyang et al. 2022; OpenAI 2023), has been utilized on downstream application tasks in different domains due to its inexpensive costs and human-aligned response (Ouyang et al. 2022). There is an increasing interest in the use of ChatGPT on social science research questions, especially questions about social networks (Ziems et al. 2023; Zhu et al. 2023). ChatGPT could be the potential solution to these two challenges. First, motivated by the superiority of ChatGPT over human annotators in various text annotation tasks (Kocoń et al. 2023), we hypothesize that ChatGPT has a deep understanding of emoji applications and can replace human workers to label emoji attributes. Second, if ChatGPT has a similar understanding of emojis to humans, when users encounter unfamiliar emojis, they can ask ChatGPT to explain the meaning of the emoji to increase information transparency. On the other hand, the previous work suggests that there exist hallucinations in LLMs’ generations (Zhang et al. 2023), inconsistencies between the model’s output and the real-world facts or user inputs. Whether ChatGPT can produce generations similar to humans for the meaning of emojis is still questionable.
@@ 3 Emoji Semantics
We start evaluating ChatGPT’s understanding on emojis by asking ChatGPT to explain the semantic meanings of the emoji. Emojis, as universal language tokens, are applied by users with different language background. In this section, we conduct a qualitative study on the LLM explanation of emoji universal semantics without any given context, and then we probe the ChatGPT understandings in the English context and Chinese context, respectively.
@@ 3.3 Emoji Semantics with Language/Cultural Context
Emojis have been used differently across languages and cultures (Lu et al. 2016; Chen et al. 2019; Wang 2022). While ChatGPT shows outstanding performances on various tasks, it also exhibits different behaviors across languages and cultures (chang2023survey,cao2023assessing). To evaluate how Emoji understanding and applications are affected by such phenomenon, we test if ChatGPT can properly explain emojis that are unique or have unique meanings in Chinese social media, i.e., [Uncaptioned image] (slightly smiling face), [Uncaptioned image] (pill), [Uncaptioned image] (broken), [Uncaptioned image] (onlooker), and [Uncaptioned image] (doge). Additionally, we include an emoji combo, [Uncaptioned image] [Uncaptioned image], that consists of a cow face and a beer mug. We prompt ChatGPT to explain the usage of these emojis. The prompts are written in English or Chinese, with or without specification of usage in Chinese pop culture. For emojis that do not have standard names (i.e., [Uncaptioned image] and [Uncaptioned image]), we prompt their pictures to GPT4V. An example is shown in Figure 3.
We find that GPT models can generate desired interpretations in all cases, but only with appropriate prompts. The prompts need to include very specific information sometimes. For instance, when prompted with “Explain [Emoji]”, GPT models failed to generate desired explanations and answered with literal interpretations from their Unicode names or visual cues (e.g., [Uncaptioned image] as friendly instead of its common ironical use in Chinese pop culture (Wang 2022)). When prompted in Chinese, only [Uncaptioned image] [Uncaptioned image] got the correct explanation. However, [Uncaptioned image] and [Uncaptioned image] get desired explanations when prompted with “Explain the [Emoji] in the context of Chinese pop culture”, no matter whether prompted in English or Chinese, while [Uncaptioned image] and [Uncaptioned image] had to use the Chinese version for GPT4V. For [Uncaptioned image], we had to specifically mention its homophonic usage with yào wán through a multi-turn interaction with ChatGPT.
@@
Meanwhile, answers in both languages are prone to hallucinations even when a large proportion of answers are desired. For example, the model falsely claims that the [Uncaptioned image] emoji was known as the “Face with Tears of Joy” emoji as shown in Figure 3. The model also linked [Uncaptioned image] to the red pill and blue pill metaphor from the Matrix movie and incorrectly states that this is a common usage in Chinese social media.
@@ 4.1 Emoji Sentiment Annotation
We first ask ChatGPT to label the 50 most frequently used emojis in Section 3.1 with positive, neutral, and negative sentiment and compare the differences between human and ChatGPT annotations. Sentiments in emojis can be determined from two aspects: visual features and real use, and the sentiments from two aspects may not be unified. For example, [Uncaptioned image] (pleading face) is negative in visual features due to tears but may behave as positive in real use. For ChatGPT, In addition to the function of LLMs (GPT4), it develops a vision language model (VLM) GPT4v and supports image input, demonstrating significant progress in image understanding (Yang et al. 2023; Wang et al. 2024). We probe GPT4v and GPT4 by providing the image containing the emojis or the Unicode to infer the emoji sentiment from visual features and from real use, respectively. We present the prompt details in Table 7(a) in Appendix. For GPT4v, to eliminate the influence of actual usage knowledge from LLMs, our prompt additionally asks VLM to only based on the visual features in the image to infer the sentiment. For the 50 selected emojis, only [Uncaptioned image] (pleading face) and [Uncaptioned image] (eyes) have different annotations for GPT4 and GPT4v. For the [Uncaptioned image] emoji, VLM attributes a negative sentiment based on its visual features: a frown, wide eyes, and raised eyebrows. Conversely, GPT4 associates the [Uncaptioned image] emoji with positive sentiment, perceiving it as an expression of vulnerability, empathy, and affection. Regarding the [Uncaptioned image] emoji, GPT4 assigns a neutral sentiment, indicative of observation, whereas GPT4v leans towards a neutral yet slightly positive sentiment, reflecting an underlying connotation of interest. These findings suggest that while GPT4v and GPT4’s sentiment annotations are generally similar, the divergent interpretations of the [Uncaptioned image] emoji may highlight the evolution of the emoji sentiment derived from the user application setting.
We also compare the GPT4 annotation with the human annotation in Hu et al. (2017), and the only difference is the sentiment annotation of emoji [Uncaptioned image] flushed face. The human annotation is neutral and the response of ChatGPT (both GPT4 and GPT4v) is negative, considering feelings of embarrassment or a sense of being overwhelmed in [Uncaptioned image]. The same sentiment annotation indicates that ChatGPT can fully understand the sentiment in the emojis. From the response from ChatGPT, it often mentions that it requires the text context to determine the emoji sentiment, so next, we will discuss the ChatGPT performance on sentiment analysis on tweets with emojis.
@@ 4.2 Sentiment Prediction with Emojis
Next, we remove the emojis from the input text and ask ChatGPT to assign a sentiment label again, and the performance accuracy decreases to 69.56% and 70.19% for GPT4 and GPT3.5, respectively. However, the accuracy decrease cannot demonstrate that ChatGPT utilizes emoji information for emoji prediction, since the accuracy change is not significant and the emoji removal may also change the ground truth sentiment for the text. We compare the sentences in which ChatGPT changes the sentiment prediction after removing the emojis and find that 157 / 146 out of 805 tweets flip the prediction labels for GPT3.5 / GPT4 annotations. We present 3 randomly selected tweets with flipped predictions for GPT4 in Table 1.
@@ 5.1 Emoji Intention Annotation without Context
We first annotate emojis without any context with seven types of candidate intentions: expressing sentiment, strengthening expression, adjusting tone, expressing humor, expressing irony, expressing intimacy, and describing content (intention 1 to intention 7), and detailed definitions can be found in Hu et al. (2017). We select 15 emojis (5 positive, 5 neutral, and 5 negative emojis) in Hu et al. (2017) and ask ChatGPT to rate the willingness to use the emoji for each intention on a 7-point scale (7 = totally willing, 1 = not willing at all), same as the annotation method in Hu et al. (2017) and the complete prompt in Table 9.
We calculate that the average difference values between the willingness scores of ChatGPT and humans from intention 1 to intention 7 are -0.13, -0.29, -0.07, -0.54, -1.49, -0.05, and 1.11, respectively, indicating that the overall intention scores for ChatGPT and human annotators are similar. From the difference in willingness scores, we also find that the understanding discrepancy between humans and ChatGPT about intentions concentrates on intentions of expressing irony and describing content. Specifically, ChatGPT assigns more scores on expressing irony and fewer scores on describing content than human annotators.
@@ 5.2 Intention of Emoji Usage in GitHub Posts
We have presented similarity to understanding the intention of using emojis for humans and ChatGPT without giving any text context. Next, we ask whether this observation can be generalized on other platforms. We evaluate the annotation of ChatGPT on the intention of emojis in the GitHub post. We collect the same 2,000 emoji posts on GitHub with human annotations in Lu et al. (2018). The candidate intention of emojis in GitHub posts includes: sentiment expressed, sentiment strengthened, tone adjusted, content described, content organized, content emphasized, non-communication use and symbol, and the details of the definitions are shown in Lu et al. (2018). We compose the prompt by adding the instruction and the definition details (the complete prompt in Table 10) and ask ChatGPT to annotate the intention of emoji usage in the specific GitHub post context. The annotated accuracy between GPT3.5 and GPT4 with the human annotation is 38. 8% and 49. 0%, respectively, indicating that GPT4 has a more consistent understanding with humans of the intention of emojis in GitHub posts. To analyze the discrepancy of understanding between humans and ChatGPT, we conduct a case study and present three randomly selected posts with different annotations in Table 2.
@@ 6 Emoji Usage Associated with Communities
Previous work has explored that emoji usage can be different when associated with different communities, such as platforms, languages, genders, and hashtags (Lu et al. 2018, 2016; Chen et al. 2018; Zhou and Ai 2022; Zhou et al. 2023a). In this part, we probe whether the knowledge of the associated emoji usage patterns is encoded in ChatGPT when given the prior information about the users, platform, or co-occurred hashtags. We compose the prompt by directly asking for the most used emojis given different community information, and ChatGPT outputs the most used emojis and its justification. The probing prompt for ChatGPT is designed as follows: Could you output 10 emojis that are associated with the {specific community}? where {specific community} is the associated community, such as female users, French users, hashtags, and the GitHub platform. We present the answer from ChatGPT and the findings from previous work in Tables 3, 4 and 5. In particular, emojis associated with French, female, and male users are explored in Lu et al. (2016); Chen et al. (2018) and emojis co-occurred with different hashtags and the GitHub platform are reflected in Zhou and Ai (2022); Zhou et al. (2023a).
User
Lu et al. (2016)
and Chen et al. (2018)
Emojis from ChatGPT
@@
Male user [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Table 3: Outputs from ChatGPT and findings from Lu et al. (2016) and Chen et al. (2018) about the emojis associated with French users, female users and male users.
Hashtag Zhou and Ai (2022) ChatGPT
@@
Table 4: Outputs from ChatGPT and findings from Zhou and Ai (2022) about the emojis associated with different hashtags.
Source Associated Emojis
Zhou et al. (2023a) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Table 5: Outputs from ChatGPT and findings from Zhou et al. (2023a) about the emojis associated with GitHub.
From Table 3, 4 and 5, we observe that ChatGPT depends on stereotypical associations with the attribute of the user or the platform to select the emojis and in most cases, stereotypical selection can extract the same emojis or emojis with semantics similar to the emojis shown in previous works. Both ChatGPT and existing literature identify certain emojis as indicative of gender preferences, such as [Uncaptioned image] (soccer) for male users and [Uncaptioned image] (cherry blossom) for female users. In the context of the GitHub platform, a comparison between prior findings and ChatGPT’s responses reveals a significant overlap, with seven emojis common to both lists. Furthermore, additional emojis identified by ChatGPT align with the distinctive characteristics of the GitHub environment. For instance, [Uncaptioned image] (penguin) is associated with Linux and [Uncaptioned image] (snake) is linked to Python programming.
@@ 8.2 Recommendation with User Identity
9 Limitations and Implications
@@
Our paper reveals ChatGPT’s precise understanding of emoji functionality and intention, and the capability could benefit further researchers and social media users. For researchers engaged in future emoji studies, leveraging human-aligned explanations of emoji semantics and intentions could significantly reduce the need for extensive crowd-sourcing efforts, particularly in multilingual contexts. This approach would enable more effective and efficient research into emoji usage and interpretation. As for social media users, they could utilize ChatGPT to decipher the semantics of unfamiliar emojis, thereby enhancing the clarity and transparency of online communication.

### 16 Feb 2024 v2

@@ 1 Introduction
Large language models (LLMs), pre-training on large corpuses, have demonstrated the remarkable achievements in multiple natural language processing tasks. Among various LLMs, ChatGPT (GPT3.5 and GPT4) (Ouyang et al. 2022; OpenAI 2023), has been utilized on downstream application tasks in different domains due to its inexpensive costs and human-aligned response (Ouyang et al. 2022). There is an increasing interest in the use of ChatGPT on social science research questions, especially questions about social networks (Ziems et al. 2023; Zhu et al. 2023). ChatGPT could be the potential solution to these two challenges. First, motivated by the superiority of ChatGPT over human annotators in various text annotation tasks (Kocoń et al. 2023), we hypothesize that ChatGPT has a deep understanding of emoji applications and can replace human workers to label emoji attributes. Second, if ChatGPT has a similar understanding of emojis to humans, when users encounter unfamiliar emojis, they can ask ChatGPT to explain the meaning of the emoji to increase information transparency. On the other hand, previous work suggests that there exist hallucinations in LLMs’ generations (Zhang et al. 2023), inconsistencies between the model’s output and the real-world facts or user inputs. Whether ChatGPT can produce generations similar to humans for the meaning of emojis is still questionable.
@@ 3 Emoji Semantics
We start evaluating ChatGPT’s understanding on emojis by asking ChatGPT to explain the semantic meanings of the emoji. Emojis, as universal language tokens, are applied by users with different language and cultural backgrounds. In this section, we conduct a qualitative study on the LLM explanation of emoji universal semantics without any given context, and then we probe the ChatGPT understandings in two different language and culture settings: the English context and the Chinese context.
@@ 3.3 Emoji Semantics with Language/Cultural Context
In addition to the application in the English context, emojis also show prevalence in other specific language and culture contexts (Chen et al. 2019), and explaining emoji semantics in the Chinese context is more challenging. With the similar experiment setup in Section 3.2, we first choose emojis that have unique meanings on Chinese social networks and test if ChatGPT can properly explain their semantics, that is, [Uncaptioned image] (slightly smiling face), [Uncaptioned image] (pill), and [Uncaptioned image] (doge). Additionally, we include an emoji combo, [Uncaptioned image] [Uncaptioned image], that consists of a cow face and a beer mug. We prompt ChatGPT to explain the usage of these emojis. The prompts are written in English or Chinese, with or without specification of usage in Chinese pop culture.
We find that GPT models can generate desired interpretations in all cases, but only with appropriate prompts. The prompts need to include very specific information sometimes. For instance, when prompted with “Explain [Emoji]”, GPT models failed to generate desired explanations and answered with literal interpretations from their Unicode names or visual cues (e.g., [Uncaptioned image] as friendly instead of its common ironical use in Chinese pop culture (Wang 2022)). When prompted in Chinese, only [Uncaptioned image] [Uncaptioned image] got the correct explanation. However, [Uncaptioned image] and [Uncaptioned image] get desired explanations when prompted with “Explain the [Emoji] in the context of Chinese pop culture”, no matter whether prompted in English or Chinese. For [Uncaptioned image], we had to specifically mention its homophonic usage with yào wán through a multi-turn interaction with ChatGPT.
@@
Meanwhile, answers in both languages are prone to hallucinations even when a large proportion of answers are desired. The model links [Uncaptioned image] to the red pill and blue pill metaphor from the Matrix movie and incorrectly states that this is a common usage in Chinese social media.
@@ 4.1 Emoji Sentiment Annotation
We first ask ChatGPT to label the 50 most frequently used emojis in Section 3.1 with positive, neutral, and negative sentiment and compare the differences between human and ChatGPT annotations. We present the prompt details in Table 8(a) in Appendix.
We compare the GPT4 annotation with the human annotation in Hu et al. (2017), and the only difference is the sentiment annotation of emoji [Uncaptioned image] flushed face. The human annotation is neutral and the response of ChatGPT is negative, considering feelings of embarrassment or a sense of being overwhelmed in [Uncaptioned image]. The same sentiment annotation indicates that ChatGPT can fully understand the sentiment in the emojis. From the response from ChatGPT, it often mentions that it requires the text context to determine the emoji sentiment, so we discuss the influence of emojis for ChatGPT on the downstream sentiment analysis task.
@@ 4.2 Sentiment Prediction with Emojis
Next, we remove the emojis from the input text and ask ChatGPT to assign a sentiment label again, and the performance accuracy decreases to 69.56% and 70.19% for GPT4 and GPT3.5, respectively. However, the accuracy decrease cannot demonstrate that ChatGPT utilizes emoji information for emoji prediction, since the accuracy change is not significant and the emoji removal may also change the ground truth sentiment for the text. We compare the sentences in which ChatGPT changes the sentiment prediction after removing the emojis and find that 19.50% / 18.13% tweets flip the prediction labels for GPT3.5 / GPT4 annotations. We present three randomly selected tweets with flipped predictions for GPT4 in Table 1.
@@ 5.1 Emoji Intention Annotation without Context
We first annotate emojis without any context to probe the understanding of ChatGPT on the universal intention. We select the same seven types of candidate intentions: expressing sentiment, strengthening expression, adjusting tone, expressing humor, expressing irony, expressing intimacy, and describing content, with previous work, and detailed definitions can be found in Hu et al. (2017). We select 15 emojis (5 positive, 5 neutral, and 5 negative emojis) in Hu et al. (2017) and ask ChatGPT to rate the willingness to use the emoji for each intention on a 7-point scale (7 = totally willing, 1 = not willing at all), same as the annotation method in Hu et al. (2017) and the complete prompt in Table 10.
We calculate that the average difference values between ChatGPT-rated scores and humans are -0.13 (expressing sentiment), -0.29 (strengthening expression), -0.07 (adjusting tone), -0.54 (expressing humor), -1.49 (expressing irony), -0.05 (expressing intimacy), and 1.11 (describing content), respectively. The average standard deviation (SD) of human-rated scores of seven intentions from Hu et al. (2017) is 1.643. ChatGPT and human annotators reach a high consensus on emojis in the intention of expressing sentiment, strengthening expressing, and adjusting tones. On other intentions, such as describing content, ChatGPT vary from human annotated scores but the score discrepancy is still smaller than the average human SD. Hence, we claim that the understanding of ChatGPT on the universal usage intention of emojis is generally similar to human users.
@@ 5.2 Intention of Emoji Usage in GitHub Posts
The similar intention annotation without context between ChatGPT and humans is consistent with the accurate explanation of emoji semantics and the precise annotation of emoji sentiment. When using emojis on different platforms, the intention could also vary. For example, users from GitHub may apply emojis to organize the content in the post, but seldom apply emojis to express irony. It is more challenging for ChatGPT to understand the emoji intention when given GitHub post context. In this part, we collect the same 2,000 emoji posts on GitHub with human annotations in Lu et al. (2018) and compare the annotation of emoji intention from ChatGPT and human ground truth labels. The candidate intention of emojis in GitHub posts is shown in Table 2, and the details of the definitions are shown in Lu et al. (2018).
Intention
Sentiment
expressed
Sentiment
strengthened
Tone
adjusted
Content
described
Accuracy 67.43% 30.85% 48.64% 47.54%
Intention
Content
organized
Content
emphasized
Non
communication use
Symbol
Accuracy 40.63% 55.49% 34.62% 14.58%
Table 2: Accuracy of ChatGPT annotation for each intention on emojis in GitHub posts with human annotation as the ground truth labels.
The annotated accuracy between GPT3.5 and GPT4 with human annotation is 38. 8% and 49. 0%, respectively. The detailed accuracy of GPT4 for each intention is presented in Table 2. We observe that for emojis to express sentiment and emphasize content in GitHub posts, ChatGPT gives more similar judgement with the human annotators. However, for symbol (symbolize emojis) and non-communication use (unintentional use), which is rarely applied in other social platforms, ChatGPT cannot generate the same intention annotation with annotators. To analyze the discrepancy of understanding between humans and ChatGPT, we conduct a case study and present three randomly selected posts with different annotations in Table 3.
@@ 6 Emoji Usage Associated with Communities
Previous work has explored that emoji usage can be different when associated with different communities, such as platforms, languages, genders, and hashtags (Lu et al. 2018, 2016; Chen et al. 2018; Zhou and Ai 2022; Zhou et al. 2023a). In this part, we probe whether the knowledge of the associated emoji usage patterns is encoded in ChatGPT when given the prior information about the users, platform, or co-occurred hashtags. We compose the prompt by directly asking for the most used emojis given different community information, and ChatGPT outputs the most used emojis and its justification. The probing prompt for ChatGPT is designed as follows: Could you output 10 emojis that are associated with the {specific community}? where {specific community} is the associated community, such as female users, French users, hashtags, and the GitHub platform. We present the answer from ChatGPT and the findings from previous work in Tables 4, 5 and 6. In particular, emojis associated with French, female, and male users are explored in Lu et al. (2016); Chen et al. (2018) and emojis co-occurred with different hashtags and the GitHub platform are reflected in Zhou and Ai (2022); Zhou et al. (2023a).
User Emoji from actual usage Emoji from ChatGPT
@@
Male user
(Chen et al. 2018)
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Table 4: Outputs from ChatGPT and emojis from actual usage (Lu et al. 2016; Chen et al. 2018) associated with French users, female users, and male users.
Hashtag Emoji from actual usage Emoji from ChatGPT
@@
Table 5: Outputs from ChatGPT and emojis from actual usage (Zhou and Ai 2022) associated with different hashtags.
Source Emojis associated with GitHub
Actual usage [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Table 6: Outputs from ChatGPT and emojis from actual usage (Zhou et al. 2023a) associated with the GitHub platform.
From Table 4, 5 and 6, we observe that ChatGPT depends on stereotypical associations with the attribute of the user or the platform to select the emojis and in most cases, stereotypical selection can extract the same emojis or emojis with semantics similar to the emojis shown in previous works. Both ChatGPT and existing literature identify certain emojis as indicative of gender preferences, such as [Uncaptioned image] (soccer) for male users and [Uncaptioned image] (cherry blossom) for female users. In the context of the GitHub platform, a comparison between prior findings and ChatGPT’s responses reveals a significant overlap, with seven emojis common to both lists. Furthermore, additional emojis identified by ChatGPT align with the distinctive characteristics of the GitHub environment. For instance, [Uncaptioned image] (penguin) is associated with Linux and [Uncaptioned image] (snake) is linked to Python programming.
@@ 8.2 Recommendation with User Identity
9 Implications and Limitations
Our paper reveals ChatGPT’s precise understanding of emoji functionality and intention. The capability of ChatGPT on emojis could benefit further researchers and social media users. For researchers engaged in future emoji studies, leveraging human-aligned explanations of emoji semantics and intentions could significantly reduce the need for extensive crowd-sourcing efforts, particularly in multilingual contexts. This approach would enable more effective and efficient research into emoji usage and interpretation. As for social media users, they could utilize ChatGPT to decipher the semantics of unfamiliar emojis, thereby enhancing the clarity and transparency of online communication.
All the experiments in our paper treat emojis as text tokens in the Unicode form. But some emojis due to its application in the specific cultural context or new appearance are not included in the Unicode organization, such as [Uncaptioned image] (broken) and [Uncaptioned image] (onlooker) in Chinese social media. For these emojis without standard Unicode, we can utilize the vision language model (VLM) from ChatGPT: GPT4V, which has been demonstrated with significant progress in image understanding (Yang et al. 2023; Wang et al. 2024), to help us understand the emoji functions. We prompt the images of [Uncaptioned image] and [Uncaptioned image] to GPT4V for semantic explanations. An example is shown in Figure 3. We observe similar patterns as in Section 3.3. GPT4V generates hallucinations with Chinese prompts but outputs the correct explanation when supplementing the prompt with the context of Chinese pop culture.
We also repeat the annotation of the sentiment of emojis in Section 3.1 on GPT4V. For the 50 selected emojis, only [Uncaptioned image] (pleading face) and [Uncaptioned image] (eyes) have different annotations for GPT4 and GPT4v. For the [Uncaptioned image] emoji, VLM attributes a negative sentiment based on its visual features: frown, wide eyes, and raised eyebrows. In contrast, GPT4 associates the [Uncaptioned image] emoji with positive sentiment, perceiving it as an expression of vulnerability, empathy, and affection. Regarding the [Uncaptioned image] emoji, GPT4 assigns a neutral sentiment, indicative of observation, while GPT4v leans towards a neutral, yet slightly positive, reflecting an underlying connotation of interest.
These findings suggest that, while the GPT4v and GPT4 sentiment annotations are generally the same, divergent interpretations of the [Uncaptioned image] emoji can highlight the evolution of the emoji sentiment derived from the user application setting. The results in GPT4V imply that, when encountering specific emojis without the standard Unicode, the notable capability in emoji understanding of GPT4V can also help researchers and users.